Find the range, the standard deviation, and the variance for the given samples. Round noninteger to the nearest tenth.

1. 48, 91, 87, 93, 59, 68, 92, 100, 81
2. 93, 67, 49, 55, 92, 87, 77, 66, 73, 96, 54
3. 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4,
4. 8, 6, 8, 6, 8, 6, 8, 6, 8, 6, 8, 6, 8
5. -8, -5, -12, -1, 4, 7, 11
6. -23, -17, -19, -5, -4, -11, -31
1
Expert's answer
2021-07-14T10:09:41-0400

1.


"48, 59, 68, 81,87,91, 92, 93, 100"




"Range=100-48=52"


"mean=\bar{x}=\dfrac{\displaystyle\sum_{i=1}^nx_i}{n}=\dfrac{1}{11}(48+59+68+81+87"

"+91+92+93+100)=\dfrac{719}{9}"

"\approx79.9""Variance=s^2=\dfrac{\displaystyle\sum_{i=1}^n(x_i-\bar{x})^2}{n-1}"

"=\dfrac{1}{8}((48-\dfrac{719}{9})^2+(59-\dfrac{719}{9})^2+(68-\dfrac{719}{9})^2"

"+(81-\dfrac{719}{9})^2+(87-\dfrac{719}{9})^2+(91-\dfrac{719}{9})^2"

"+(92-\dfrac{719}{9})^2+(93-\dfrac{719}{9})^2+(100-\dfrac{719}{9})^2)"

"\approx311.6"

"s=\sqrt{s^2}\approx17.7"


2,


"49,54,55,66,67,73, 77, 87, 92,93, 96"




"Range=96-49=47""mean=\bar{x}=\dfrac{\displaystyle\sum_{i=1}^nx_i}{n}=\dfrac{1}{11}(49+54+55+66+67"

"+73+77+87+92+93+96)=\dfrac{809}{11}"

"\approx73.5""Variance=s^2=\dfrac{\displaystyle\sum_{i=1}^n(x_i-\bar{x})^2}{n-1}"

"=\dfrac{1}{10}((49-\dfrac{809}{11})^2+(54-\dfrac{809}{11})^2+(55-\dfrac{809}{11})^2"

"+(66-\dfrac{809}{11})^2+(67-\dfrac{809}{11})^2+(73-\dfrac{809}{11})^2"

"+(77-\dfrac{809}{11})^2+(87-\dfrac{809}{11})^2+(92-\dfrac{809}{11})^2"

"+(93-\dfrac{809}{11})^2+(96-\dfrac{809}{11})^2)"

"\approx284.5"

"s=\sqrt{s^2}\approx16.9"


3.



"4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4"




"Range=4-4=0"


"mean=\bar{x}=\dfrac{\displaystyle\sum_{i=1}^nx_i}{n}=\dfrac{1}{17}(4+4+4+4+4"

"+4+4+4+4+4+4+4+4+4+4+4+4)"

"=4""Variance=s^2=\dfrac{\displaystyle\sum_{i=1}^n(x_i-\bar{x})^2}{n-1}"

"=\dfrac{1}{16}((4-4)^2+(4-4)^2+(4-4)^2+(4-4)^2"

"+(4-4)^2+(4-4)^2+(4-4)^2+(4-4)^2+(4-4)^2"

"+(4-4)^2+(4-4)^2+(4-4)^2+(4-4)^2+"

"+(4-4)^2+(4-4)^2+(4-4)^2+(4-4)^2)=0"

"s=\sqrt{s^2}=0"

4.


"6, 6, 6, 6, 6, 6, 8, 8, 8, 8, 8, 8, 8"




"Range=8-6=2"


"mean=\bar{x}=\dfrac{\displaystyle\sum_{i=1}^nx_i}{n}=\dfrac{1}{13}(6+6+6+6+6+6"

"+8+8+8+8+8+8+8)=\dfrac{92}{13}"

"\approx7.1""Variance=s^2=\dfrac{\displaystyle\sum_{i=1}^n(x_i-\bar{x})^2}{n-1}"

"=\dfrac{1}{12}((6-\dfrac{92}{13})^2+(6-\dfrac{92}{13})^2+(6-\dfrac{92}{13})^2"

"+(6-\dfrac{92}{13})^2+(6-\dfrac{92}{13})^2+(6-\dfrac{92}{13})^2"

"+(8-\dfrac{92}{13})^2+(8-\dfrac{92}{13})^2+(8-\dfrac{92}{13})^2"

"+(8-\dfrac{92}{13})^2+(8-\dfrac{92}{13})^2+(8-\dfrac{92}{13})^2"

"+(8-\dfrac{92}{13})^2)\approx1.1"

"s=\sqrt{s^2}\approx1.0"

5.


"-12,-8, -5,-1, 4, 7, 11"




"Range=11-(-12)=23""mean=\bar{x}=\dfrac{\displaystyle\sum_{i=1}^nx_i}{n}=\dfrac{1}{7}(4-12-8-5-1+4"

"+7+11)=\dfrac{-4}{7}\approx-0.6"

"Variance=s^2=\dfrac{\displaystyle\sum_{i=1}^n(x_i-\bar{x})^2}{n-1}"

"=\dfrac{1}{6}((-12-\dfrac{-4}{7})^2+(-8-\dfrac{-4}{7})^2+(-5-\dfrac{-4}{7})^2"

"+(-1-\dfrac{-4}{7})^2+(4-\dfrac{-4}{7})^2+(7-\dfrac{-4}{7})^2"

"+(11-\dfrac{-4}{7})^2)\approx69.6"

"s=\sqrt{s^2}\approx8.3"



6.


"-31,-23, -19, -17, -11,-5, -4"




"Range=-4-(-31)=27""mean=\bar{x}=\dfrac{\displaystyle\sum_{i=1}^nx_i}{n}=\dfrac{1}{7}(-31-23-19-17"

"-11-5-4)=\dfrac{-110}{7}\approx-15.7"

"Variance=s^2=\dfrac{\displaystyle\sum_{i=1}^n(x_i-\bar{x})^2}{n-1}"

"=\dfrac{1}{6}((-31-\dfrac{-110}{7})^2+(-23-\dfrac{-110}{7})^2"

"(-19-\dfrac{-110}{7})^2+(-17-\dfrac{-110}{7})^2"

"(-11-\dfrac{-110}{7})^2+(-5-\dfrac{-110}{7})^2"

"+(-4-\dfrac{-110}{7})^2\approx95.6"

"s=\sqrt{s^2}\approx9.8"